OpenStudy (smurfy14):

[2x-1]/[(x+2)(x^2+3)]=0 solve the equation.

OpenStudy (anonymous):

x=1/2

OpenStudy (smurfy14):

how did you get it?

OpenStudy (anonymous):

hold on let me type it out

OpenStudy (smurfy14):

ok

OpenStudy (anonymous):

hold on im going to type it out on word then ill attach it, its just easier

OpenStudy (smurfy14):

alright

myininaya (myininaya):

the bottom doesn't make a fraction 0 only the top does so set 2x-1=0 and solve for x 2x=1 x=1/2 but you also have to be careful that this doesn't make the bottom zero since 0/0 does not exsit but you are good the answer is 1/2

OpenStudy (smurfy14):

oh ok but how come you dont do anything to the bottom?

OpenStudy (anonymous):

I have a much longer description lol click on the file i have attached

myininaya (myininaya):

the only thing that makes a fraction zero is the top

OpenStudy (smurfy14):

oh ok it makes sense thanks guys :)

OpenStudy (smurfy14):

do either of yall no how to do this one? x^3-2x^2-5x+6=0

OpenStudy (anonymous):

x=1,3,-2